The length of one leg of a right triangle is twice the length of
the other. The hypotenuse is 10 cm. Find the lengths of the legs.

Answer by rfer(10417)

(2x)^2+(x)^2=10^2
5x^2=100
x^2=20
x=4.47
2x=8.94